OpenStudy (anonymous):

Find a nonzero matrix A for which A^2=0.

OpenStudy (anonymous):

@Hero, @thomaster

OpenStudy (loser66):

try \[\left[\begin{matrix}1-i~~&1+i\\1+i~~&1-i\end{matrix}\right]\]

OpenStudy (loser66):

@zepdrix can you help her a real entries matrix?

OpenStudy (loser66):

for real entries matrix, you have a lot of them. just take (A-\(\lambda I)^2\) you have 0 matrix. for example: \(A = \left[\begin{matrix}0&-2\\2&4\end{matrix}\right]\) so, eigenvalue is 2 double roots. Then \(A-\lambda I= \left[\begin{matrix}-2&-2\\2&2\end{matrix}\right]\) . This matrix square = 0 matrix. ALWAYS. From this, you can construct any matrix whose square =0